OpenStudy (anonymous):

What is the derivative of x^(x^(1/2))?

OpenStudy (anonymous):

I put it in Wolfram Alpha and this is what they gave me, but I have no idea how they got it: 1/2x^(x^(1/2)-1/2)(log(x) + 2)

OpenStudy (anonymous):

is it equal to something. I want to use ln. If you times it by the natural log you can bring the power out in front and use the product rule to differentiate

OpenStudy (anonymous):

thats probably what they did

OpenStudy (anonymous):

so. (x^1/2)( ln x)

OpenStudy (anonymous):

The actual problem is: Find the limit of that equation as x goes to 0+.

OpenStudy (anonymous):

What calc is this

OpenStudy (anonymous):

Calculus 2. This is an indeterminate form problem

OpenStudy (anonymous):

You end up getting 0^(0) which requires L'hospital's rule

OpenStudy (anonymous):

Which requires taking the derivative of that :(

OpenStudy (anonymous):

The limit appears to go to 1.

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

I can find the derivative but honestly I can't remember L'hospital's rule. I vaguely remember taking notes on this years ago. Hopefully somebody else can help. I don't want to steer you in the wrong direction

OpenStudy (anonymous):

I'm not sure L'Hopital's would help here since it's not a ratio. It's not like a 0/0 or ∞/∞.

OpenStudy (anonymous):

Cliffsedge, L'hopital's rule is done when you get: 00, 0/0, 1∞, ∞ − ∞, ∞/∞, 0 × ∞, and ∞0. (from Wikipedia)

OpenStudy (anonymous):

So in this case, you get 0^0 which is one of the known ones :/.

OpenStudy (anonymous):

Understood, but what is the process of L'Hopital's Rule?

OpenStudy (anonymous):

L'Hopital's rule says that if you get in a situation like this, you take the derivative of it. F(x)/g(x) becomes f'(x) / g'(x)

OpenStudy (anonymous):

For mine, since we don't have a fraction, we just take the derivative, and see if plugging 0 again would give us a situation like the ones I listed. If I do, I use it again, if I don't, then whatever I get is my final answer.

OpenStudy (anonymous):

If I keep getting loops, and keep getting stuck, and have no other choices to do anything, then it's an undefined problem.

OpenStudy (anonymous):

I don't think you can just take the derivative right off like that. You can exponentiate first, then move the limit inside the exponent.

OpenStudy (anonymous):

hmm, the professor does very simple examples in class -.-

OpenStudy (anonymous):

\[x^{\sqrt{x}} \rightarrow e^{\ln(x^{\sqrt{x}})} \rightarrow e^{\sqrt{x}\ln(x)} \]

OpenStudy (anonymous):

hmm I think that's right, but it requires another L'hopital's rule, because I just plugged in 0, and it doesn't work

OpenStudy (anonymous):

Yes, now you can find \[\lim_{x \rightarrow 0+}\sqrt{x}\ln(x) = \lim_{x \rightarrow 0+}\frac{\ln(x)}{x^{-1/2}}\]

OpenStudy (anonymous):

Now regular L'Hopital's rule will break that one down and you should see that it goes to zero.

OpenStudy (anonymous):

What happened to the e?

OpenStudy (anonymous):

It's still there. I just moved the limit evaluation to the exponent to show that it is e^0 =1.

OpenStudy (anonymous):

hmm now if I plug in 0, it still does the same thing.

OpenStudy (anonymous):

If you still get 0/0, use L'Hopital's again until the denominator goes to 1.

OpenStudy (anonymous):

the bottom becomes -1/2x^(1/2)

OpenStudy (anonymous):

Check your power rule again.

OpenStudy (anonymous):

Hmm. x^(-1/2) = -1/2x^(1/2) * (1), no?

OpenStudy (anonymous):

subtract one..ahh I was adding lol

OpenStudy (anonymous):

I have been doing integrals all day long, so I just confused myself even more. -1/2x^(-3/2)

OpenStudy (anonymous):

\[\frac{d}{dx}x^{-1/2} = -\frac{1}{2}x^{-3/2}\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I feel 'ya. That happens to me all the time. :-p

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

Just keep simplifying and differentiating top and bottom until you get -3x^(1/2) over 1.

OpenStudy (anonymous):

we have never used l'hopital's rule this many times o.o. But okay

OpenStudy (anonymous):

Yeah, usually it does its thing in one or two applications, but sometimes you have to keep putting those functions through the wringer until they come out nice and smooth the way you want them.

OpenStudy (anonymous):

okay, I will keep doing it. Thank you so much :)

OpenStudy (anonymous):

You're welcome. This was a nasty one. Similar kind of annoyance as having to do integration-by-parts over and over again.

OpenStudy (anonymous):

I think that's why I didn't understand it, integration by parts is actually next chapter's material..and the professor gave us some questions that require it, although we don't know it yet lol.

OpenStudy (anonymous):

Some professors think you need to learn to fly before you learn to walk . . .